# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ B.3 Proof of the Preliminary Results
Taking conditional expectation on both sides in (49) yields
@@
Recall that $0<\lambda\leq\min\{\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]/(4\mathbb{E}[\sigma_{\tau}^{4}m_{\tau}^{4}]),1/(2\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}])\}$. Using Proposition B.1 and the stochastic gradient (14), we have
@@
Proof of Lemma B.4.
Applying Itô’s formula and using the process (9) with $s$ given in (13), we have, for any $t\in[0,T]$,
@@
Integrating both side and using Corollary B.3, we obtain
@@
Proof of Lemma B.5.
Using the process (10) with the approximating function $s$ given in (13), we have
@@
Taking conditional expectation on both sides in (50), using the independence of $Y_{k}^{\text{EM}}$ and $\bar{Z}_{k+1}$, Young’s inequality and $20<\gamma\leq 1/$ 2, we obtain
@@
Using Corollary B.3 in (51) yields
@@
Proof of Lemma B.6.
Using the process (11) with the approximating function $s$ given in (13), Lemma B.5, Corollary B.3 and $\gamma\in(0,1/2]$, we have, for any $t\in[0,T]$,
@@
Proof of Lemma B.7.
Applying Itô’s formula to the process (11) with the approximating function $s$ given in (13), we have, for any $t\in[0,T]$,
@@
Using Lemma B.6, we have
@@
Integrating both sides, using $\gamma\in(0,1/2]$ and Corollary B.3 yields
@@
Appendix C Proofs of the Results in the General Case
In this section, we provide the proof of Theorem 3.10. We start by introducing the results which will be used in the proof of Theorem 3.10.
C.1 Preliminary Estimates for the General Case
Throughout this section, we fix $\epsilon\in(0,1)$. The following auxiliary results will be used in the proof of Theorem 3.10 and their proofs are postponed to Appendix C.3.
We provide an upper bound for the moments of $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ defined in (11).
Lemma C.1.
Let Assumptions 1 and 3 hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
@@ C.1 Preliminary Estimates for the General Case
The following result provides an estimate for the one step error associated with $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ defined in (11).
Lemma C.2.
Let Assumptions 1 and 3 hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
@@
The following result is a modification of (Kumar and Sabanis, 2019, Lemma 4.1).
Lemma C.3.
Let Assumption 3 hold and let $b:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ such that
@@
Lemma C.4.
Let Assumption 3 hold and let $b$ be as in (52). Then, one obtains that, for any $(t,\theta,x)\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{M}$ and $k=1,\dots M$,
@@
C.2 Proof of the Main Result for the General Case
Proof of Theorem 3.10.
We proceed as in the proof of Theorem 3.1 using the splitting
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))+W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))\\ &\quad+W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))+W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}})),\end{split}$ (54)
where the first term on the right-hand side in (54) corresponds to the error made by the early stopping and the remaining terms have the same interpretation of the corresponding ones in (29).
Upper bound on $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}})).$
@@ C.2 Proof of the Main Result for the General Case
Recall that $t_{K}=T-\epsilon$. Using the representation of the OU process (2) and the inequalities (55), we have
$\begin{split}W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))&=W_{2}(\pi_{\mathsf{D}},\mathscr{L}(X_{T-t_{K}}))\\ &\leq\sqrt{\mathbb{E}\left[|X_{0}-X_{T-t_{K}}|^{2}\right]}\\ &\leq\sqrt{2}\left[(1-m_{T-t_{K}})\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sigma_{T-t_{K}}\sqrt{M}\right]\\ &\leq\sqrt{2}\sigma_{T-t_{K}}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})\\ &\leq 2\sqrt{\epsilon}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M}).\end{split}$ (56)
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}})).$
@@
By integrating, taking expectations on both sides in (57), and using Remark 3.4 with the lower bound $\widehat{L}_{\text{MO}}$ in the estimate (19), we have
@@
Using (59), we have
$\begin{split}W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}]}\\ &\leq\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}.\end{split}$ (60)
Upper bound on $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}})).$
@@
By integrating and taking the expectation on both sides in (61), and using the Remark 3.4 with the lower bound $\widehat{L}_{\text{MO}}$ in the estimate (19), Young’s inequality with $\zeta\in(0,1)$ and Assumption 4, we have
@@
Using (62) and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y^{\text{aux}}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{K}}-Y_{t_{K}}^{\text{aux}}|^{2}]}\\ &\leq\sqrt{2\zeta^{-1}}e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)}\sqrt{\varepsilon_{\text{SN}}}.\end{split}$ (63)
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}})).$
The following bound is derived by modifying (Kumar and Sabanis, 2019, Lemma 4.7). For the sake of presentation, let $b:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ such that
@@
Consequently, $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ can be expressed using (64) as
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=b(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}})\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{M}).$ (65)
@@
Integrating and taking the expectation on both sides in (66), using Cauchy–Schwarz inequality, Young’s inequality with $\zeta\in(0,1)$, Assumption 3 and Remark 3.3, yield
@@
We proceed estimating the third term on the right-hand side of (67). Using (65), Young’s inequality with $\zeta\in(0,1)$, Lemma C.3 and Lemma C.2, yields, for any $t\in[0,T-\epsilon]$
@@
$\begin{split}&\int_{0}^{t}\mathbb{E}\left[\langle Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}},b(T-s,\hat{\theta},\widehat{Y}_{s}^{\text{EM}})-\ b(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})\rangle\right]\mathrm{d}s\\ &=\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\Bigg{(}b^{(k)}(T-s,\hat{\theta},\widehat{Y}_{s}^{\text{EM}})-b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad-\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}(\widehat{Y}^{\text{EM},(j)}_{s}-\widehat{Y}^{\text{EM},(j)}_{\lfloor s/\gamma\rfloor\gamma})\Bigg{)}\ \mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}(\widehat{Y}^{\text{EM},(j)}_{s}-\widehat{Y}^{\text{EM},(j)}_{\lfloor s/\gamma\rfloor\gamma})\right)\ \mathrm{d}s\right]\end{split}$

### v5

@@ C.3 Proof of the Preliminary Results
Taking conditional expectation on both sides in 54 yields
@@
Recall that $0<\lambda\leq\min\{\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]/(4\mathbb{E}[\sigma_{\tau}^{4}m_{\tau}^{4}]),1/(2\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}])\}$. Using Proposition 13 and the stochastic gradient 14, we have
@@
Proof of Lemma 16.
Applying Itô’s formula and using the process 9 with $s$ given in 13, we have, for any $t\in[0,T]$,
@@
Integrating both side and using Corollary 15, we obtain
@@
Proof of Lemma 17.
Using the process 10 with the approximating function $s$ given in 13, we have
@@
Taking conditional expectation on both sides in 55, using the independence of $Y_{k}^{\text{EM}}$ and $\bar{Z}_{k+1}$, Young’s inequality and $20<\gamma\leq 1/$ 2, we obtain
@@
Using Corollary 15 in 56 yields
@@
Proof of Lemma 18.
Using the process 11 with the approximating function $s$ given in 13, Lemma 17, Corollary 15 and $\gamma\in(0,1/2]$, we have, for any $t\in[0,T]$,
@@
Proof of Lemma 19.
Applying Itô’s formula to the process 11 with the approximating function $s$ given in 13, we have, for any $t\in[0,T]$,
@@
Using Lemma 18, we have
@@
Integrating both sides, using $\gamma\in(0,1/2]$ and Corollary 15 yields
@@
Appendix D Proofs of the Results in the General Case
In this section, we provide the proof of Theorem 10. We start by introducing the results which will be used in the proof of Theorem 10.
D.1 Preliminary Estimates for the General Case
Throughout this section, we fix $\epsilon\in(0,1)$. The following auxiliary results will be used in the proof of Theorem 10 and their proofs are postponed to Appendix D.3.
We provide an upper bound for the moments of $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ defined in 11.
Lemma 20.
Let Assumptions 1 and 3.b hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
@@ D.1 Preliminary Estimates for the General Case
The following result provides an estimate for the one step error associated with $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ defined in 11.
Lemma 21.
Let Assumptions 1 and 3.b hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
@@
The following result is a modification of Kumar & Sabanis (2019, Lemma 4.1).
Lemma 22.
Let Assumption 3.b hold and let $b:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ such that
@@
Lemma 23.
Let Assumption 3.b hold and let $b$ be as in 57. Then, one obtains that, for any $(t,\theta,x)\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{M}$ and $k=1,\dots M$,
@@
D.2 Proof of the Main Result for the General Case
Proof of Theorem 10.
We proceed as in the proof of Theorem 1 using the splitting
$\begin{split}W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{K}}))+W_{2}(\mathcal{L}(Y_{t_{K}}),\mathcal{L}(\widetilde{Y}_{t_{K}}))\\ &\quad+W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K}}),\mathcal{L}(Y_{t_{K}}^{\text{aux}}))+W_{2}(\mathcal{L}(Y_{t_{K}}^{\text{aux}}),\mathcal{L}(Y_{K}^{\text{EM}})),\end{split}$ (59)
where the first term on the right-hand side in 59 corresponds to the error made by the early stopping and the remaining terms have the same interpretation of the corresponding ones in 34.
Upper bound on $W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{K}})).$
@@ D.2 Proof of the Main Result for the General Case
Recall that $t_{K}=T-\epsilon$. Using the representation of the OU process 2 and the inequalities 60, we have
$\begin{split}W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{K}}))&=W_{2}(\pi_{\mathsf{D}},\mathcal{L}(X_{T-t_{K}}))\\ &\leq\sqrt{\mathbb{E}\left[|X_{0}-X_{T-t_{K}}|^{2}\right]}\\ &\leq\sqrt{2}\left[(1-m_{T-t_{K}})\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sigma_{T-t_{K}}\sqrt{M}\right]\\ &\leq\sqrt{2}\sigma_{T-t_{K}}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})\\ &\leq 2\sqrt{\epsilon}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M}).\end{split}$ (61)
Upper bound on $W_{2}(\mathcal{L}(Y_{t_{K}}),\mathcal{L}(\widetilde{Y}_{t_{K}})).$
@@
By integrating, taking expectations on both sides in 62, and using Remark 4 with the lower bound $\widehat{L}_{\text{MO}}$ in the estimate 19, we have
@@
Using 64, we have
$\begin{split}W_{2}(\mathcal{L}(Y_{t_{K}}),\mathcal{L}(\widetilde{Y}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}]}\\ &\leq\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}.\end{split}$ (65)
Upper bound on $W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K}}),\mathcal{L}(Y_{t_{K}}^{\text{aux}})).$
@@
By integrating and taking the expectation on both sides in 66, and using the Remark 4 with the lower bound $\widehat{L}_{\text{MO}}$ in the estimate 19, Young’s inequality with $\zeta\in(0,1)$ and Assumption 4, we have
@@
Using 67 and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K}}),\mathcal{L}(Y^{\text{aux}}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{K}}-Y_{t_{K}}^{\text{aux}}|^{2}]}\\ &\leq\sqrt{2\zeta^{-1}}e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)}\sqrt{\varepsilon_{\text{SN}}}.\end{split}$ (68)
Upper bound on $W_{2}(\mathcal{L}(Y_{t_{K}}^{\text{aux}}),\mathcal{L}(Y_{K}^{\text{EM}})).$
The following bound is derived by modifying Kumar & Sabanis (2019, Lemma 4.7). For the sake of presentation, let $b:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ such that
@@
Consequently, $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ can be expressed using 69 as
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=b(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}})\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathcal{N}(0,I_{M}).$ (70)
@@
Integrating and taking the expectation on both sides in 71, using Cauchy–Schwarz inequality, Young’s inequality with $\zeta\in(0,1)$, Assumption 3.b and Remark 3, yield
@@
We proceed estimating the third term on the right-hand side of 72. Using 70, Young’s inequality with $\zeta\in(0,1)$, Lemma 22 and Lemma 21, yields, for any $t\in[0,T-\epsilon]$
@@
$\begin{split}&\int_{0}^{t}\mathbb{E}\left[\langle Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}},b(T-s,\hat{\theta},\widehat{Y}_{s}^{\text{EM}})-\ b(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})\rangle\right]\mathrm{d}s\\ &=\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\Bigg{(}b^{(k)}(T-s,\hat{\theta},\widehat{Y}_{s}^{\text{EM}})-b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad-\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}(\widehat{Y}^{\text{EM},(j)}_{s}-\widehat{Y}^{\text{EM},(j)}_{\lfloor s/\gamma\rfloor\gamma})\Bigg{)}\ \mathrm{d}s\Bigg{]}\\ &\quad+\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\left(Y_{s}^{\text{aux},(k)}-\widehat{Y}_{s}^{\text{EM},(k)}\right)\left(\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}(\widehat{Y}^{\text{EM},(j)}_{s}-\widehat{Y}^{\text{EM},(j)}_{\lfloor s/\gamma\rfloor\gamma})\right)\ \mathrm{d}s\right]\\ &\leq\frac{\zeta}{2}\int_{0}^{t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\ \mathrm{d}s\\ &\quad+\frac{1}{2\zeta}\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}|b^{(k)}(T-s,\hat{\theta},\widehat{Y}_{s}^{\text{EM}})-b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})\end{split}$